OpenStudy (anonymous):

The population of a town increases at the rate of 5 % every year.The present population is 8000.In how many years it will become 9200?

OpenStudy (anonymous):

please help me

OpenStudy (kenljw):

8000 (1+.05)^y = 9200 1.05^y=9.2/8 y log(1.05)=log(9.2/8) y=2.86

OpenStudy (anonymous):

(1+.05) from where it comes?

OpenStudy (anonymous):

ok i got it ,thanks

OpenStudy (kenljw):

start 8000 first year 8000(1+.05) second year 8000(1+.05)(1+.05)=8000(1+.05)^2 General 8000(1+.05)^y the 1+.05 represents the growth per year